A person is at 25 meters and ends at 2 meters and the trip took 4.5 seconds. What was the average speed
1
Expert's answer
2020-09-21T06:29:15-0400

Distance travelled is "=25-2=23 m"

Time taken is "=4.5 s"

The average speed is given by

"v_{avg}=\frac{Total distance}{Total time taken}"

"=\frac{23 m}{4.5 s}=5.1 m/s"